The length of a rectangle is five units less than its width,w. Write an expression that represents the area of the rectangle.
Answer by JulietG(1812)

Usually, the length is the larger number, but...
L = W -5
Length * Width = Area
Substitute the known value of L into the equation.
(W-5) * W = Area
w^2 - 5W = area
.
So, for instance, if the width is 10, then
L = 10-5
L = 5
5 * 10 = 50 (area)
which is
W^2 - 5W
10^2 - 5(10)
100 - 50 = 50